OpenStudy (cheska_p):

First derivative of sec^2(x)+tan^2(x). The answer is 4sec^2(x)tanx, but I'm not sure how. Could anyone give my a step my step?

myininaya (myininaya):

one way: you could use Pythagorean identity then differentiate

myininaya (myininaya):

recall 1+tan^2(x)=sec^2(x)

OpenStudy (cheska_p):

Let me try and work that out. I'll let you know if I get it this time

OpenStudy (cheska_p):

@myininaya I did get the correct answer this time. Thank you! Just wondering, would there be another way to get that answer without using the pythagorean trig identity?

myininaya (myininaya):

Let's see what happens if we differentiate the initial thing you have there without applying any identities \[2 \sec(x) \cdot \sec(x) \tan(x)+2\tan(x) \cdot \sec^2(x) \\\] we don't need any identities

OpenStudy (dinamix):

\[(\sec^2x)'=2\sec^2xtanx\] \[\tan^2x=2\tan(x)\sec^2x \] cuz (tan(x))'=sec^2(x)

myininaya (myininaya):

these are like terms

OpenStudy (dinamix):

this my method @myininaya @cheska_P

OpenStudy (dinamix):

\[=\tan(x)[2\sec^2x+2\sec^2x]\]

OpenStudy (cheska_p):

@dinamix I'm so happy you did that method. I was doing that way originally and not getting the answer, BUT I completely forgot the derivative of tanx gives you sec^2x. I kept writing secx and didn't understand why I wasn't getting the answer! Thank you!!

OpenStudy (dinamix):

@cheska_P u are welcome